An RLC series circuit with 150-ohm resistor, 25-mH inductor, and 2 microfarad-capacitor is
powered by an AC voltage source with a peak voltage of 340V and a frequency of 660 Hz. (a)
Find the peak current that flows in the circuit. (b) Determine the phase angle of the source
voltage relative to the current. (c) Determine the peak voltage across the resistor and its phase
angle relative to the voltage source. (d) Find the peak voltage across the inductor and its phase
angle relative to the voltage source. (e) Find the peak voltage across the capacitor and its phase
angle relative to the voltage source.
Answer #1

Inductive Reactance, X_L = 2 \pi f L = 2 \pi * 660 * 25 * 10^{-3} =103.67 \Omega

Capactive Reactance, X_C = \frac{1}{2 \pi f C} = \frac{1}{2 \pi * 660 * 2* 10^{-6}} =120.57 \Omega

Impedance, Z = \sqrt{R^2 + (X_L - X_C)^2 } = \sqrt{150^2 + (103.67 - 120.57)^2 } = 150.95 \Omega

a)

Peak Current, I = \frac{340}{150.95 } = 2.25 A \rightarrow Answer

b)

Phase Angle, \phi = \tan^{-1}\left [ \frac{X_L - X_C}{R} \right ] = \tan^{-1}\left [ \frac{103.67- 120.57}{150} \right ] = -6.43^o \rightarrow Answer

c)

Peak voltage across the resistor, V_R = I R = 2.25 * 150 = 337.5 V

Phase angle relative to the voltage source is Zero

d)

Peak voltage across the Inductor, V_L = I X_L = 2.25 * 103.67 = 233.26 V

Phase angle relative to the voltage source, \phi = \tan^{-1}\left [ \frac{V_L}{V_R} \right ] = \tan^{-1}\left [ \frac{233.26}{337.5} \right ] = 34.5^o

e)

Peak voltage across the Capacitor, V_L = I X_C = 2.25 * 120.57 = 271.28 V

Phase angle relative to the voltage source, \phi = \tan^{-1}\left [ \frac{V_C}{V_R} \right ] = \tan^{-1}\left [ \frac{-271.28}{337.5} \right ] = -38.8^o
Answer #2

To get the peak current, we must find the impedence 


image.png

answered by: anonymous